### v1

# pnasresearcharticle

Abstract
Decades after being replaced with digital platforms, analogue computing has experienced a surging interest following developments in metamaterials and intricate fabrication techniques. Specifically, wave-based analogue computers which impart spatial transformations on an incident wavefront, commensurate with a desired mathematical operation, have gained traction owing to their ability to directly encode the input in its unprocessed form, bypassing analogue-to-digital conversion. While promising, these systems are inherently limited to single-task configurations. Their inability to concurrently perform multiple tasks, or compute in parallel, represents a major hindrance to advancing conceptual mechanical devices with broader computational capabilities. In here, we present a first attempt to simultaneously process independent computational tasks within the same architected structure. By breaking time invariance in a set of metasurface building blocks, multiple frequency-shifted beams are self-generated which absorb notable energy amounts from the fundamental signal. The onset of these tunable harmonics, enables distinct computational tasks to be assigned to different independent “channels”, effectively allowing an analogue mechanical computer to multitask.
keywords:
mechanical computing |||| metamaterial |||| acoustics |||| wave scattering
18]2
[…]
Mechanical computing refers to a form of computing where well-established mechanical mechanisms such as levers or fluidics carry out an intended operation (1). This form, which relies on both analogue (e.g., sliders, curved surfaces) and discrete mechanisms (e.g., gears, pinwheels), historically predates digital computing, and was used to conduct astronomical calculations (2), estimate tide heights (3), and even model the economy (4), long before digital computers appeared on the scene (5). Analogue computers in particular are those which exploit a continuous variation in a physical parameter (e.g., voltage or mechanical deformation) to perform a computation (6). In the modern era, they were effectively replaced by their digital counterparts which addressed the limitations of analogue computers, providing compact, noise-free and and high-speed computations. Over the past decade, however, analogue computing research regained impetus owing to recent advances in smart materials and engineered structures (i.e., metamaterials), combined with additive manufacturing and novel fabrication techniques (7, 8). Today’s generation of mechanical computers (MCs) propose novel ways to process bit abstraction and mechanical logic (9), and span a wide portfolio of mechanisms, including rotary joints (10), conductive polymers (11), bistable lattices (12), origami systems (13, 14), and various micro-mechanical elements (15).
Tuned wave scattering in dispersive media, especially with the advent of engineered materials and waveguides, provides a rich platform for analogue computing (16). Understandably, the spurt of activity in wave-based computational metamaterials originated in photonic media, owing to ultrafast propagation speeds, potentially allowing analogue photonic computers to outpace digital electronics (17, 18, 19, 20, 21, 22, 23). In addition to operating at subwavelength scales due to small optical wavelengths, the analogue nature of these computers enables them to bypass analogue-to-digital conversion requirements which tend to consume significant energy (24). On the other hand, despite the inherently lower speeds of elastoacoustic waves, the value of integrating computational (25) and neuromorphic (26) functionality within a self-contained mechanical system has triggered a stream of new paradigms in the domain of wave-based analogue mechanical computers (AMCs), i.e., those which employ targetted scattering of elastic and acoustic waves to carry out a prescribed mathematical operation (27, 28). In remote or low access conditions, the ability to sustain a bare minimum of computational readiness is invaluable, especially for inputs which already exist in mechanical or vibroacoustic waveform. Example circumstances include elevated temperatures which are detrimental to electronic parts, ionizing radiation which trigger leakage currents, and off-grid environments where digital computing is not an option (29, 30, 31).
The emphasis of this work is on a class of AMCs which undertakes complex high-order computations such as differentiation, integration, and convolution by manipulating impinging mechanical waves (32). One methodology is the Green’s function metamaterial approach (33, 34), where the Green’s function of the operator of choice is directly realized in real space using, for instance, phase-shifted Bragg grating (35, 36), and high-index metamaterials (28). Alternatively, the metasurface (MS) approach (19, 37, 38, 39) adopted here allows the realization of a broader range of complex operators by implementing subwavelength metasurfaces positioned between spatial Fourier transformers, foregoing the need to integrate sophisticated geometries (7). The aforementioned approaches, while effective, are inherently limited to single-task configurations, as exemplified by Fig. 1A. As such, to conduct multiple operations, current state-of-the-art AMCs can either execute two identical tasks (e.g., spatial differentiation of two different incident signals) one after the other, i.e., in series, or two distinct tasks (e.g., spatial differentiation and integration of an incident signal) via two independent AMCs, as depicted in Fig. 1B. This inability to concurrently perform multiple tasks represents a major obstacle to advancing mechanical computing, and crossing this hurdle can be conducive to conceptual devices with broader computational capabilities (40).
In this work, we present a first successful attempt to concurrently process independent operations within the same architected structure, thus effectively unlocking parallel computing in AMCs. To accomplish this, we will tap into the dynamical features of time-variant periodic media (41), building off nonreciprocal wave phenomena and the manipulation of monochromatic incident waves to induce supplementary waves with spectral footprints that are distinct from the primary one (42, 43). By exploiting time-modulated metasurfaces as building blocks of an AMC, we will demonstrate the system’s ability to instigate multiple scattered wave beams, representing frequency conversions (harmonics) of a single input, that simultaneously propagate within distinct channels along the AMC’s operational spectrum. Since these frequencies can be independently tuned as functions of the constitutive parameters of the unit cell geometry and physical properties, they can be potentially assigned distinct computational tasks, thus allowing parallel operations to take place. As a result, upon completion, the computing output can be directly extracted from the culminating frequency-domain waveforms at the AMC’s readout plane at the respective frequencies, as graphically shown in Fig. 1C. The system is based on acoustic metasurface layers structured from the well-established Helmholtz resonator unit cells (44, 45, 46). We will show that the configuration of such resonators effectively alters the characteristics of propagating waves, enabling the realization of metasurfaces that can impose the desired transmission and phase profiles and subsequently impart various operations on an incoming wave.
Theory
Metasurface approach
The analogue mechanical computer (AMC), depicted in Fig. 2A, consists of three main components: a spatial Fourier transform sub-block (FT), an operator metasurface or a space-filtering sub-block (SF), and an inverse Fourier transform sub-block (IFT), adopting the well-developed metasurface approach (7) wherein a mathematical operation is applied to an input function $f(y)$, spatially encoded in the form of an incident wave, and transforming it to a corresponding output $g(y)$, via the following scheme:
$g(y)=\textrm{IFT}\big{[}H(k_{y})\cdot\textrm{FT}[f(y)]\big{]}$ (1)
where $k_{y}$ is the spatial frequency. In such an AMC, the input wave field is introduced to the system at a given frequency, henceforth denoted as the fundamental operational frequency. As the wave propagates through the first sub-block, which is in principle a focusing metasurface (47, 48), it exhibits a Fourier transformation by which the input function gets transferred to the spatial Fourier (frequency) domain. The focusing metasurface makes use of Snell’s law and its transmission coefficient $\bar{T}_{\mathrm{FT}}$ is designed such that:
$\bar{T}_{\mathrm{FT}}=e^{\mathrm{i}\frac{2\pi}{\lambda}\sqrt{y^{2}+\ell_{f}^{2}}}$ (2)
where $\mathrm{i}$ is the imaginary unit, $\lambda$ is the wavelength associated with the input wave frequency, and $\ell_{f}$ is the intended focal length. With $\bar{T}_{\mathrm{FT}}$ having an absolute value of unity, it can be inferred from the equation that a near-full transmission is required for all $50$ unit cells constituting the focusing metasurface, while satisfying the required phase profile illustrated in Fig. 2C. The latter is realized by minimizing $|\phi(y)-\phi_{\mathrm{FT}}(y)|$ via utilizing the available parameters shown in the design maps (Fig. 2B), with $\phi$ and $\phi_{\mathrm{FT}}$ being the actual and target phase angles applied at the focusing metasurface unit cells. It is worth noting that the focusing takes effect over the distance from the center of the focusing metasurface to the center of the subsequent sub-block, thus requiring $\ell_{f}=\ell+w$ for waves to be effectively focused at the intended location, where $\ell$ and $w$ represent the dimensions defined in Fig. 2A. Although the current design employs a focal length exceeding several wavelengths (see SI Appendix, section S1), the integration of subwavelength focusing techniques offers substantial size reduction capabilities, promoting a compact design (49, 50, 51).
The FT operation sets up the wave for the mathematical computation which is conducted by the following sub-block, the operator metasurface. By rendering the input function in the frequency domain, calculus-type operations can be carried out via simple algebraic multipliers. This is realized by tailoring the operator metasurface’s transmission profile to satisfy the transfer function $H(k_{y})$ associated with the operation of interest. For instance, the transmission coefficients required to obtain a first derivative and an integration of the input function are:
$\bar{T}_{\mathrm{diff}}=\mathrm{i}\Big{(}\frac{2y}{w}\Big{)}$ (3)
and,
$\bar{T}_{\mathrm{int}}=-\mathrm{i}\Big{(}\frac{d}{y}\Big{)}$ (4)
respectively, where $d<<w$ is an arbitrary normalizing length (16, 38). Similar to the focusing layer, the $50$ unit cells of the operator metasurface are carefully selected from the design maps, with the goal of minimizing the deviation of the actual transmission amplitude $|\bar{T}(y)|$ from the target one $|\bar{T}_{\mathrm{SF}}(y)|$, which corresponds to the operation-specific transmission function, e.g., $\bar{T}_{\mathrm{diff}}$ and $\bar{T}_{\mathrm{int}}$ for spatial differentiation and integration, respectively. We note that while the focusing layer’s transmission profile is symmetric, the operator layer’s profile is not. This is attributed to the fact that two different phase angles shifted by $\pi$ are required for the metasurface’s two halves ($y\in[-\frac{w}{2},0]$ and $y\in[0,\frac{w}{2}]$), as shown in Fig. 2C. Finally, the wave undergoes another Fourier transformation at the IFT sub-block, which reverts it back to the time domain, and the desired output wave is received at the system’s final terminal, i.e., the mechanical computer’s readout plane.
Unit cell
The dynamics governing the AMC necessitate a unit cell with a wide range of both transmission amplitudes $|\bar{T}|$ and phase angles $\phi$. A subwavelength unit cell is adopted for that purpose which consists of a straight pipe coupled with four shunted Helmholtz resonators (HRs), as schematically illustrated in Fig. 2A (52, 53). To achieve near-full transmission through the unit cell for the focusing metasurfaces (FT and IFT sub-blocks), the straight pipe is constructed as half-wavelength long ($w=\lambda/2$) to match the acoustic impedance of the incident waves. Additionally, the combination of four HRs acting as lumped elements, generates the necessary effective acoustic reactance, enabling the full $2\pi$ phase range to be achieved by tuning the height $h_{3}$. However, the majority of mathematical operations in the SF sub-block require transmission profiles with varied amplitudes, typically ranging from 00 to $1$, along with the corresponding full $2\pi$ phase range. By incorporating the height of the straight pipe, $h_{1}$, as a tunable parameter, we limit the flow partially, allowing for precise control of the transmission amplitude. While the fine tuning of $h_{1,3}$ allows the unit cell to satisfy the aforementioned design requirements, its structural simplicity enables straightforward implementation, facilitating both simulations (SI Appendix, Fig. S2) and experimental realizations (54).
Utilizing this unit cell configuration to construct the three AMC components, two separate AMCs are designed to perform the individual mathematical operations of spatial differentiation and integration following the procedure explained earlier. Without loss of generality, we use a Gaussian function, $p_{i}(y)=ye^{-100y^{2}}$, as an input function at a frequency of $3.45$ kHz. The latter is equivalent to the resonant frequency of the unit cell’s straight pipe, yielding a wavelength of $\approx 0.1\leavevmode\nobreak\ \mathrm{m}$ for atmospheric air. Note that $p_{i}(y)$ represents shape of the input waveform and that the ensuing process is amplitude-independent. As such, it can be scaled as required for any given application (e.g., amplitude of an incident pressure wave). The results, depicted in SI Appendix, Fig. S3, show the system’s ability to accurately execute both operations, and serve as a benchmark for the rest of the study.
Temporal modulation
The metasurface approach, outlined thus far, is neither capable of carrying out more than one operation at once, nor performing a different operation than the one it is specifically configured for (e.g., differentiation or integration). As such, a single AMC can only undertake two tasks in series (Fig. 1B), and both tasks have to be identical (e.g., two differentiations or two integrations depending on what the SF block is designed to accomplish). This inability to concurrently perform multiple tasks represents a major hindrance in current state-of-the-art AMCs, further exacerbating the problem of low computational speeds. To unlock parallel computing in AMCs, we exploit targetted changes in the frequency content of a propagating signal associated with periodic media which carry a momentum bias, and stemming from the temporal modulation of its unit cell properties. Across photonic (55), acoustic (56, 57, 58), and elastic phased structures (59, 60, 61), these modulations have been shown to redistribute notable portions of the bulk wave energy into side bands representing up and down-converted harmonics of the fundamental frequency (Fig. 3A). To enable these frequency conversions, the height of the shunted resonators in the AMC’s unit cell is modulated at a rate, $\omega_{m}=0.25$ kHz, that is much smaller than the fundamental $3.45$ kHz frequency to ensure a stable response while still allowing wave amplitudes at different frequency channels to be controllable. The modulation frequency is kept the same for all unit cells for simplicity. The selected time modulation is not only practically feasible in a number of ways, e.g., by pumping fluid into/out of the unit cell (54), but is also computationally inexpensive to realize.
In our numerical finite element model, each resonator domain is defined as a moving mesh where deformed mesh positions are introduced as degrees of freedom in the system dynamics (SI Appendix, section S4). The top boundaries of the resonators are set to have a reciprocating vertical motion, $A_{m}\cos{(\omega_{m}t+\phi_{m})}$, where $A_{m}$ and $\phi_{m}$ indicate the modulation amplitude and phase, respectively. Side boundaries are set to have zero normal displacement, ensuring the mesh deformation occurs in the desired direction. Parametric studies are then conducted with varying tunable heights $h_{1}$ and $h_{3}$, and modulation parameters $A_{m}$ and $\phi_{m}$. Specifically, for every configuration, a time-dependent simulation is run with the unit cell embedded within a waveguide structure (SI Appendix, Fig. S5), and a subsequent analysis of the simulation data using a Fast Fourier Transform (FFT) provides discrete values for the transmission amplitudes and phases at select frequencies (See Fig. 3B).
Results and Discussion
Single computing
As a sanity check, and for the sake of future comparison, we first introduce a uniform time modulation to a system that is originally designed to conduct a single mathematical operation, namely differentiation, meaning that the modulation parameters $A_{m}$ and $\phi_{m}$ are fixed for all unit cells in the operator MS. Despite the anticipated energy redistribution from the fundamental frequency channel to other harmonics, the AMC’s functionality is expected to remain consistent across all frequency channels due to the uniformity of the imposed modulation. A wavefield snapshot of this computer, depicted in Fig. 3C, illustrates the system’s ability to extract the spatial derivative of the input load, i.e., $dp_{i}(y)/dy$, at the readout plane. Additionally, the rightmost panel of the figure confirms the ability of the AMC to successfully compute $dp_{i}(y)/dy$ at both the fundamental ($3.45$ kHz) and down-converted ($3.20$ kHz) frequency channels, albeit with a slightly lower amplitude for the latter.
Parallel computing
Following this confirmation, the AMC is now redesigned to concurrently run two independent operations in parallel, a capability that has been thus far elusive. The modulation parameters as well as the tunable heights are carefully selected such that the highest pressure amplitudes take place at the fundamental ($\omega=3.45$ kHz) and down-converted ($\omega-\omega_{m}=3.20$ kHz) frequencies, as depicted by the color-coded bars in Fig. 3B. The goal is for the AMC to fulfil a primary mathematical operation (differentiation) at the fundamental frequency of the input signal, while simultaneously running a secondary operation (integration) encrypted within a “hidden” layer known to the user, which corresponds to the down-converted frequency.
[…]
Insensitivity to task type
We place emphasis on the fact that this newly-embedded AMC ability to conduct computations in parallel is highly robust. In other words, the system maintains its functionality regardless of the required operation or the frequency channel in which it is implemented. Figure 4 shows the aforementioned AMC emulating the performance displayed in Fig. 3, but with differentiation and integration being performed at the flipped channels, i.e., the down-converted ($\omega-\omega_{m}=3.20$ kHz) and fundamental ($\omega=3.45$ kHz) frequencies, respectively. Despite substantial variations between the two transmission profiles leading to entirely different designs of the tuned parameters, the system maintained its functionality, as depicted in the figure. The results confirms the AMC’s insensitivity to specific operation profiles and a greater dependence on the available unit cell designs and their inherent performance. In addition to the previously discussed advantages, the FFT of the design unit cells, shown in Fig. 3B, shows notable transmission amplitudes at several other harmonics along the operational spectrum, indicating the potential to perform more than merely two parallel operations. This capability can be effectively exploited by optimizing the modulation frequency as well as the tuning parameters of the underlying metasurfaces.
In summary, this work has demonstrated the successful realization of parallel computing within the realm of analogue mechanical computers that exploit targetted, guided wave scattering in tuned metasurfaces. The presented system was shown to be capable of concurrently executing two distinct mathematical operations, by breaking off monochromatic incident waves into supplementary signals, two of which were carefully tailored to output the intended computations on the input signal. The process was enabled by the utilization of acoustic metasurface unit cells composed of Helmholtz resonating cavities, the geometrical parameters of which exhibiting finely tuned temporal modulations. Notably, we have shown this approach to be robust and effectively insensitive to the type of computation or the frequency in which it is hosted. The inherent multi-tasking capability of the shown metamaterial-based system presents exciting avenues for future exploration in the domains of physical and reservoir computing, both in mechanical media and beyond.
[…]
This work was supported by the US Army Research Office (ARO) under award no. W911NF-23-1-0078 as well as the US National Science Foundation under award no. 1847254.
[…]